OpenStudy (anonymous):

~Slope~ I need to check myself with this equation! Which of the following is the slope between the two points (-3, 1) and (3, -6)? 6/7 -6/7 7/6 -7/6 Okay Here Is My Work: Point one-(x^1,y^1)=(-3,1) Point two-(x^2,y^2)=(3,6) Slope=y^2-y^1/x^2-x^1=-6-1/3--3=5/0 Okay so this is not in any of the choices so I thought about adding it instead: -6+1/3+-3=7/6 That is in the choices but I don't know whether it's a Negative or a Positive.

OpenStudy (anonymous):

m=-3/2

OpenStudy (raffle_snaffle):

^what he said

OpenStudy (anonymous):

Um...No that is not one of the choices!

OpenStudy (anonymous):

It is just a check...but if I have the wrong answer than I need to know :P

OpenStudy (anonymous):

Okay, so since 1 - (-6) = +7 that goes over -3 -3 = -6 you get 7/-6 which means that the whole slope would be negative! So... -7/6 You were so close! :)

OpenStudy (anonymous):

Thanks so much :P, I didn't know whether it was Negative or Positive :P

OpenStudy (anonymous):

See I had the answer, That's why I check; instead of asking for the answer :P

OpenStudy (anonymous):

Glad I could help! Hopefully I did! That's the best way to do it, too :)

OpenStudy (anonymous):

You did...I fanned you :D Thanks dear :P